### v1

@@ 3 Results
We evaluate our results mainly on the GliODIL dataset with 152 patients. We measure the recurrence coverage, that is, the ratio of tumor recurrence, which is covered by the proposed radiation plan. A qualitative assessment of how our method compares against the current standard procedure is shown in Figure 2. The visual assessment highlights the crucial role of the edema region. Although it is ignored in the standard procedure, there is a clear indication that edema has a predictive correlation with tumor recurrence. We find this is also the largest benefit over the best baseline method, "static grid discretization".
@@
Figure 2: Demonstration of our method on example patients. In the first row, we show the two input MR images with the tumor and the recurrence, that should be covered. Edema is shown in blue, enhancing tumor in green, and necrotic in red. Our method predicts a continuous assumption of tumor cells shown in the second row. This continuous concentration is thresholded to have the same volume as the standard plan (grey) to create the CTV (orange). In the last row, we compare our method to the standard plan for different patients. the same result for additional patients from the RHUH dataset.
@@
The quantitative results of our method are shown in Table 1. We compare our data to the state-of-the-art models in our field, [3, 7, 16]. We see that our method outperforms all existing methods in predicting any recurrence, meaning edema, necrosis, or enhancing core, and also in predicting only contrast-enhancing tumor recurrence, while only the worst thresholds lead to significant results, which occurs due to the applied non-parametric rank test.
@@
Table 1: Comparison of recurrence segmentation coverage given equal radiation volume, tested for different edema and core thresholds (Figure 3). Our method outperforms all others with short runtime.
Recurrence Coverage - GliODIL (152 Patients) Any [%] Enhancing Core [%] Runtime
NN (Unconstrained) $65.38\pm 2.05$ $69.02\pm 2.79$ < 1 min
NN (Physics-Constrained) $62.06\pm 2.11$ $75.25\pm 2.84$ < 1 min
Numerical Physics Simulations $61.16\pm 2.12$ $75.34\pm 2.87$ 2 h
Static Grid Discretization $67.80\pm 2.09$ $84.42\pm 2.40$ 30 min
Standard Plan $63.59\pm 2.26$ $82.42\pm 2.60$ < 1 min
@@
Ours Median Thresholds $70.93\pm 1.99^{\ddagger**}$ $85.02\pm 2.31$ 1 min
@@
As it is unclear which tumor concentrations are visible in MR images, we tested an extensive, medically plausible range of thresholds $\tau_{\text{Core}}$ and $\tau_{\text{Edema}}$. For our comparison on recurrence prediction, these tumor thresholds can also be interpreted as the weighting between the importance of tumor core vs. edema. We show the sweep over those thresholds in Figure 3.
Figure 3: The recurrence coverage is shown for the GliODIL Dataset with 152 patients comparing contrast-enhancing recurrence (orange) and any recurrence (green). Different core and edema visibility threshold parameters $\tau$ are tested. On the right, the individual results for each patient are shown compared to the standard plan. A clear improvement is visible for many patients, while also a lot of patients result in 0% or 100% coverage.
In general, we do not find a large variation for different thresholds. Predicting the enhancing recurrence is slightly easier when selecting a high core threshold. For the "any recurrence" prediction, we find an optimal outcome at ($\tau_{\text{Core}}=0.6$, $\tau_{\text{Edema}}=0.2$). We assume this relatively even distribution of the tumor concentration range between 0 and 1 leaves a good distance between two thresholds, holding an optimal range that fits the physics loss.
3.1 Replication Study on RHUH Dataset
To validate our findings, we compare our method with the standard plan and the best baseline method on the independent RHUH dataset, which stems from another center. The results are shown in Table 2. We find similar results to those on the primary dataset. For "any recurrence," we see a significant improvement of 5% to 8% compared to the standard plan. Identical to the GliODIL dataset, the improvement in predicting the contrast-enhancing part of the results is about 2% to 3%.
Table 2: We evaluated our method on the independent second RHUH Dataset with 40 patients. We compared our method to the standard plan and the best-performing baseline from the GliODIL dataset.
Recurrence Coverage - RHUH (40 Patients) Any [%] Enhancing Core [%]
Static Grid Discretization $70.95\pm 3.29$ $85.36\pm 5.4$
Standard Plan $65.47\pm 3.51$ $85.06\pm 5.4$
Ours Worst $70.27\pm 3.28^{**}$ $86.91\pm 4.93$
Ours Median $72.02\pm 3.21^{**}$ $87.15\pm 5.03$
Ours Best $\mathbf{73.18}\pm 3.16^{\dagger**}$ $\mathbf{87.48}\pm 4.90$
3.2 Extending Our Framework
We evaluate how our method can be extended. Therefore, we conduct two additional experiments, where we (i) include advanced imaging data and (ii) add a further physical loss into the tumor cell estimation.
3.2.1 Including PET Imaging
@@
As an example of how our method can be extended with additional measurements, we evaluate the effects of an additional PET loss. We tested 58 patients who had amino acid PET images in the GliODIL dataset. The same experiment was conducted by [2] in a recent NeurIPS publication, simulating not only the influence of PET but also advanced brain deformations. In Table 3, we compare the default method with optimal parameters and our method with the additional PET loss to this method and to the other methods also used for the full GliODIL dataset (Table 1). Both the default method and the one including PET loss show a clear improvement over the standard plan and over the dynamic grid discretization in terms of "any recurrence" prediction. In enhancing recurrence coverage, the dynamic grid discretization is better than our method, while our methods are slightly still outperforming the standard plan. For this, the distribution is also highly dominated by outliers, as most patients already have 100% enhancing recurrence coverage in the standard plan.
Table 3: Comparison of our additional method utilizing PET imaging. We use the subset of 58 patients of the Gliod dataset having PET imaging. Additionally, we can compare our results to the method "Dynamic Grid Discretization"[2], which requires PET.
Recurrence Coverage with PET (58 Patients) Any [%] Enhancing Core [%]
NN (Unconstrained) $59.0\pm 4.3$ $66.8\pm 4.9$
NN (Physics-Constrained) $70.4\pm 3.7$ $84.3\pm 3.3$
Numerical Physics Simulations $67.1\pm 3.8$ $86.2\pm 3.6$
Static Grid Discretization $72.9\pm 3.5$ $89.0\pm 3.3$
Dynamic Grid Discretization $74.7\pm 3.1$ $\mathbf{89.9}\pm 2.7$
Standard Plan $70.0\pm 3.8$ $87.3\pm 3.6$
Ours $76.2\pm 3.4^{\ddagger**}$ $88.2\pm 3.4$
Ours with additional PET Loss $\mathbf{77.4}\pm 3.3^{\ddagger**}$ $88.4\pm 3.3$
3.2.2 Including Further Physical Constraints
As an additional loss, we test how the assumption of a wave-like solution affects the recurrence prediction. We ran the same tumor visibility parameter sweep as shown in Figure 3 with the additional wave loss. The results are compared to the default loss and are shown in Table 4. We see a clear improvement for the median and the worst set of thresholds in predicting "any recurrence". Additionally, by using the wave loss, we get significant results for the prediction of the enhancing core. This can be interpreted as increased robustness towards the selection of hyperparameters, paving the way for further adjustment of advanced physics constraints and highlighting how extending our framework with additional physical constraints can improve model safety, an important prerequisite for clinical translation.
@@ 3.2.2 Including Further Physical Constraints
Table 4: Recurrence prediction with the additional assumption of a wave-like solution compared to the default version of our method. We compare the minimum, maximum, and median of recurrence coverage with multiple thresholds as shown in Figure 3.
Recurrence Coverage - with Wave Loss (152 Patients) Any [%] Enhancing Core [%]
@@
Ours Median Thresholds $70.93\pm 1.99^{\ddagger**}$ $85.02\pm 2.31$
Ours Best Thresholds $\mathbf{72.48}\pm 1.99^{\ddagger**}$ $85.19\pm 2.28$
Ours with Wave Loss Worst Thresholds $71.24\pm 2.08^{\ddagger**}$ $84.79\pm 2.34$
@@
4 Discussion
In this study, we demonstrated that our approach, which efficiently combines soft physical constraints and direct optimization of 3D tumor concentration to imaging data, shows promising results for recurrence prediction while substantially reducing computational demand compared to other methods. With a runtime of under one minute, our approach is markedly faster than the 30+ minutes typically required by current state-of-the-art methods. By eliminating the need for extensive simulations of tumor growth dynamics, our method facilitates practical integration into existing radiotherapy workflows, thereby enhancing the efficiency of patient-specific treatment planning.
4.1 Limitations
It is relevant to acknowledge that our method does not capture the complete trajectory of tumor growth, which may limit its utility in understanding past tumor growth dynamics. While some extrapolation may be possible based on the optimized tumor concentration, our approach does not support retrospective modeling. Furthermore, the application of radiation was conducted solely based on standard planning methods, which could potentially influence recurrence rates in the treated region.
4.2 Outlook
Importantly, our framework can easily be extended to incorporate advanced imaging methods as well as additional physical constraints. Our experimental results indicate that these enhancements, particularly the inclusion of extra physical constraints, can "stabilize" predictions and provide essential safety guarantees, which are critical for clinical translation and effective radiotherapy planning. Moreover, this versatility suggests that our framework could be applied to a broader range of medical challenges beyond brain tumor modeling, such as predicting the spread of infection.

### v2

@@ 3 Results
We evaluate our results mainly on the GliODIL dataset with 152 patients. A qualitative assessment of how our method compares against the current clinical gold standard is shown in Figure 2. Although it is ignored in the clinical standard, there is a clear indication that edema has a predictive correlation with tumor recurrence. This is also the most significant benefit over the best baseline method, "static grid discretization".
@@
Figure 2: a.) Demonstration of our method on example patients. In the first row, we show the two input MR images with the tumor and the recurrence that should be covered. Edema is shown in blue, enhancing tumor in green, and necrotic in red. Our method predicts a continuous estimation of tumor cells, as shown in the second row. This continuous concentration is thresholded to have the same volume as the standard plan (grey) to create the CTV (orange). In the last row, we compare our method to the standard plan for different RHUH patients. b.) The recurrence coverage is shown for the GliODIL Dataset with 152 patients. The individual results for each patient are shown compared to the standard plan. A clear improvement is visible for many patients, while also a lot of patients result in 0% or 100% coverage. c.) We are comparing any recurrence (green) and contrast-enhancing recurrence (orange) coverage improvement over the standard plan for different core and edema visibility threshold parameters $\tau$.
As it is unclear which tumor concentrations are visible in MR images, we tested an extensive, medically plausible range of thresholds $\tau_{\text{Core}}$ and $\tau_{\text{Edema}}$. For our comparison on recurrence prediction, these tumor thresholds can also be interpreted as the weighting between the importance of tumor core vs. edema. We show the sweep over those thresholds in Figure 2c. In general, we do not find a large variation for different thresholds. Predicting the enhancing recurrence is slightly easier when selecting a high core threshold. For the "any recurrence" prediction, we find an optimal outcome at ($\tau_{\text{Core}}=0.6$, $\tau_{\text{Edema}}=0.2$).
@@
The quantitative results are shown in Table 1, where we compare our data to the state-of-the-art models, [3, 7, 16]. We see that our method significantly outperforms all existing methods in predicting "any recurrence", meaning edema, necrosis, or enhancing core, and also in predicting only contrast-enhancing tumor recurrence while only partly significant.
To validate our findings, we compare our method with the standard plan and the best baseline method on the independent RHUH dataset, which stems from another center. The results are shown in Table 2a. We find similar results to those on the primary dataset. For "any recurrence", we see a significant improvement of 5% to 8% compared to the standard plan. Identical to the GliODIL dataset, the improvement in predicting the contrast-enhancing part of the results is about 2% to 3%.
@@
Table 1: Comparison of recurrence segmentation coverage given equal radiation volume, tested for different edema and core thresholds (Figure 2) on the GliODIL dataset with 152 patients. Our method outperforms all others with short runtime.
Recurrence Coverage - GliODIL Any [%] Enhancing Core [%] Runtime
NN (Unconstrained) [3] 65.38±2.0565.38$\pm$ 2.0565.38 ± 2.05 69.02±2.7969.02$\pm$ 2.7969.02 ± 2.79 < 1 min
NN (Physics-Constrained) [7] 62.06±2.1162.06$\pm$ 2.1162.06 ± 2.11 75.25±2.8475.25$\pm$ 2.8475.25 ± 2.84 < 1 min
Numerical Physics Simulations [16] 61.16±2.1261.16$\pm$ 2.1261.16 ± 2.12 75.34±2.8775.34$\pm$ 2.8775.34 ± 2.87 2 h
Static Grid Discretization [3] 67.80±2.0967.80$\pm$ 2.0967.80 ± 2.09 84.42±2.4084.42$\pm$ 2.4084.42 ± 2.40 30 min
Standard Plan 63.59±2.2663.59$\pm$ 2.2663.59 ± 2.26 82.42±2.6082.42$\pm$ 2.6082.42 ± 2.60 < 1 min
@@
Ours Median Thresholds $70.93\pm 1.99^{\ddagger**}$ 85.02±2.3185.02$\pm$ 2.3185.02 ± 2.31 1 min
@@
3.0.1 Including PET Imaging:
@@
As an example of how our method can be extended with additional measurements, we evaluate the effects of an additional PET loss. We tested 58 patients with amino acid PET images in the GliODIL dataset. The same experiment was conducted by [2] in a recent NeurIPS publication, simulating not only the influence of PET but also advanced brain deformations. In Table 2, we compare the default method with optimal parameters and our method with the additional PET loss to this method and the other methods also used for the full GliODIL dataset (Table 1). Both the default method and the one including PET loss show a clear improvement over the standard plan and the dynamic grid discretization regarding "any recurrence" prediction. In enhancing recurrence coverage, the dynamic grid discretization is better than our method, while our methods are still outperforming the clinical standard plan.
@@
Table 2: a.) We evaluated our method on the independent second RHUH Dataset with 40 patients. We compared our method to the standard plan and the best-performing baseline from the GliODIL dataset. b.) Comparison of our additional method utilizing PET imaging. We use the subset of 58 patients of the Gliod dataset having PET imaging. Additionally, we can compare our results to the method "Dynamic Grid Discretization"[2], which requires PET. c.) Recurrence prediction with the additional assumption of a wave-like solution compared to the default version of our method. We compare the minimum, maximum, and median of recurrence coverage with multiple thresholds as shown in Figure 2.
Experiment - Recurrence Coverage Any [%] Enhancing Core [%]
a.) RHUH (40 Patients)
Static Grid Discretization 70.95±3.2970.95$\pm$ 3.2970.95 ± 3.29 79.59±4.8279.59$\pm$ 4.8279.59 ± 4.82
Standard Plan 65.47±3.5165.47$\pm$ 3.5165.47 ± 3.51 80.69±5.0780.69$\pm$ 5.0780.69 ± 5.07
Ours Worst $70.27\pm 3.28^{**}$ 80.70±4.6080.70$\pm$ 4.6080.70 ± 4.60
Ours Median $72.02\pm 3.21^{**}$ 81.89±4.7381.89$\pm$ 4.7381.89 ± 4.73
Ours Best $\mathbf{73.18}\pm 3.16^{\dagger**}$ $\mathbf{82.06}\pm 4.70$
b.) With PET Images (58 Patients)
NN (Unconstrained) 59.0±4.359.0$\pm$ 4.359.0 ± 4.3 66.8±4.966.8$\pm$ 4.966.8 ± 4.9
NN (Physics-Constrained) 70.4±3.770.4$\pm$ 3.770.4 ± 3.7 84.3±3.384.3$\pm$ 3.384.3 ± 3.3
Numerical Physics Simulations 67.1±3.867.1$\pm$ 3.867.1 ± 3.8 86.2±3.686.2$\pm$ 3.686.2 ± 3.6
Static Grid Discretization 72.9±3.572.9$\pm$ 3.572.9 ± 3.5 89.0±3.389.0$\pm$ 3.389.0 ± 3.3
Dynamic Grid Discretization 74.7±3.174.7$\pm$ 3.174.7 ± 3.1 $\mathbf{89.9}\pm 2.7$
Standard Plan 70.0±3.870.0$\pm$ 3.870.0 ± 3.8 87.3±3.687.3$\pm$ 3.687.3 ± 3.6
Ours $76.2\pm 3.4^{\ddagger**}$ 88.2±3.488.2$\pm$ 3.488.2 ± 3.4
Ours with additional PET Loss $\mathbf{77.4}\pm 3.3^{\ddagger**}$ 88.4±3.388.4$\pm$ 3.388.4 ± 3.3
c.) With Wave Loss (152 Patients)
@@
Ours Median Thresholds $70.93\pm 1.99^{\ddagger**}$ 85.02±2.3185.02$\pm$ 2.3185.02 ± 2.31
Ours Best Thresholds $\mathbf{72.48}\pm 1.99^{\ddagger**}$ 85.19±2.2885.19$\pm$ 2.2885.19 ± 2.28
Ours with Wave Loss Worst Thresholds $71.24\pm 2.08^{\ddagger**}$ 84.79±2.3484.79$\pm$ 2.3484.79 ± 2.34
@@
3.0.2 Including Physical Constraints:
We also test how the assumption of a wave-like solution, i.e., adding further physical constraints to the estimated tumor cell concentration, affects the recurrence prediction. We ran the same tumor visibility parameter sweep as shown in Figure 2 with the additional wave loss. The results are compared to the default loss and are shown in Table 1. We see a clear improvement for the median and the worst set of thresholds in predicting "any recurrence". Additionally, by adding the wave loss, we obtain a significant improvement in the prediction of enhancing core recurrence. This can be interpreted as increased robustness towards the selection of hyperparameters, paving the way for further adjustment of advanced physics constraints and highlighting how extending our framework with additional physical constraints can improve model safety, an important prerequisite for clinical translation.
4 Conclusion
Our method estimates the 3D concentration of tumor cells by aligning it with MRI tumor segmentation while enforcing a smooth concentration landscape. This approach cuts runtimes from 30 minutes to under one minute and significantly improves recurrence prediction for 192 patients. This streamlined approach challenges the demand for complex simulations by demonstrating that simpler, efficient techniques can effectively address the challenging problem of tumor recurrence prediction. Our experimental results indicate that additional loss terms, particularly the inclusion of extra physical constraints, can stabilize predictions and provide essential safety guarantees, which are critical for clinical translation and effective radiotherapy planning. Also, this highlights the flexibility of our framework, which can thus be readily adapted to other pathologies and modalities.